### v1

# Highly edge-connected subgraphs

Abstract
Mader conjectured that an average degree of at least 3k13k-1 is sufficient for the existence of a $(k+1)$-connected subgraph. The following minimum degree version holds: Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices. Moreover, for triangle-free graphs, already an average degree of at least $2k$ is sufficient for a $(k+1)$-connected subgraph (which has at least $2(k+1)$ vertices).
For edge-connectedness (in simple graphs), we prove the following: Every graph of average degree at least $2k$ has a $(k+1)$-edge-connected subgraph on more than $2k$ vertices. Moreover, for every small $\alpha>0$ and for $k$ large enough in terms of $\alpha$, already a minimum degree of at least $k+k^{\frac{1}{2}+\alpha}=\big{(}1+o(1)\big{)}k$ is sufficient for a $(k+1)$-edge-connected subgraph.
It is shown that all of these results are sharp in some sense. The results can be used for the decomposition of graphs into two highly (edge-)connected parts.
Sufficient minimum degree conditions for the existence
of highly connected or edge-connected subgraphs
Maximilian Krone
Technische Universität Ilmenau, August 11, 2025
On highly (vertex-)connected subgraphs
Mader conjectured in [1] that every graph with average degree 3k13k-1 has a $(k+1)$-connected subgraph. This conjecture turned out to be quite hard, since the best known sufficient linear factor was improved in many iterations, but never reached $3$. We consider another version of the problem by demanding even a minimum degree of 3k13k-1. In exchange for the stronger requirement, the $(k+1)$-connected subgraph is enforced to have a non-trivial size of more than $2k$ vertices. This bonus outcome cannot be expected from an average degree of 3k13k-1, since by a construction of Carmesin in [2], there are graphs of average degree $\big{(}3+\frac{1}{3}\big{)}k-\mathcal{O}(k)$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
Theorem 1
Let $k\geq 1$ be fixed. Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices.
Proof.  Let $G$ be a graph with $v(G)>2k$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
Let $v\in V(G)$ be a vertex of degree $d_{G}(v)$. We consider a delimited penalty
$m_{G}(v)\leavevmode\nobreak\ :=\leavevmode\nobreak\ 0\lor\big{(}(3k-1)-d_{G}(v)\big{)}\land k\leavevmode\nobreak\ \in\{0,\dots,k\}\,.$
We prove by induction on $v(G)>2k$ that
$M(G)\leavevmode\nobreak\ :=\leavevmode\nobreak\ \sum_{v\in V(G)}m_{G}(v)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2k^{2}\,.$
Since $v(G)>2k$, $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$, that is, two induced but not spanning subgraphs $A$ and $B$ with $V(A)\cup V(B)=V(G)$,  $E(A)\cup E(B)=E(G)$ and $|V(A)\cap V(B)|=k$. By symmetry, we may assume that $v(A)\geq v(B)$.
Let $a\in V(A)\setminus V(B)$, $b\in V(B)\setminus V(A)$ and $s\in V(A)\cap V(B)$. We have $d_{G}(a)=d_{A}(a)$ and hence $m_{G}(a)=m_{A}(a)$. Similarly, $m_{G}(b)=m_{B}(b)$.
Case I:  $v(A)\leq 2k$ and $v(B)\leq 2k$.
We have $d_{G}(s)\leq v(G)-1\leq 3k-1$, and hence $m_{G}(s)\geq\big{(}(3k-1)-(v(G)-1)\big{)}\land k=3k-v(G)$. It is $m_{G}(a)=m_{G}(b)=k$. Hence
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{s}\big{(}3k-v(G)\big{)}+\sum_{a,b}k\leavevmode\nobreak\ \geq\leavevmode\nobreak\ k(3k-v(G))+(v(G)-k)k\leavevmode\nobreak\ =\leavevmode\nobreak\ 2k^{2}\,.$
Case II:  $v(A)>2k$ and $v(B)\leq 2k$.
We use the induction hypothesis for $A$. We have $d_{G}(s)-d_{A}(s)\leq v(B)-k$. Since the function $x\mapsto 0\lor\big{(}(3k-1)-x\big{)}\land k$ is non-expanding, this yields $m_{A}(s)-m_{G}(s)\leq v(B)-k$. Again, $m_{G}(b)=k$. Hence
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{a}m_{A}(a)+\sum_{s}\big{(}m_{A}(a)-(v(B)-k)\big{)}+\sum_{b}k$
$=\leavevmode\nobreak\ M(A)-k(v(B)-k)+(v(B)-k)k\leavevmode\nobreak\ =\leavevmode\nobreak\ M(A)\leavevmode\nobreak\ \geq\leavevmode\nobreak 2k^{2}\,.$
Case III:  $v(A)>2k$ and $v(B)>2k$.
We use the induction hypothesis for $A$ and $B$. We have $m_{G}(s)\geq 0\geq m_{A}(s)+m_{B}(s)-2k$. Hence
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{a}m_{A}(a)+\sum_{s}\big{(}m_{A}(s)+m_{B}(s)-2k\big{)}+\sum_{b}m_{B}(b)$
$=\leavevmode\nobreak\ M(A)+M(B)-k(2k)\leavevmode\nobreak\ \geq\leavevmode\nobreak 2k^{2}+2k^{2}-2k^{2}\leavevmode\nobreak\ =\leavevmode\nobreak 2k^{2}\,.$
This finishes the proof: If $G$ has minimum degree at least 3k13k-1, then $v(G)>2k$ and $M(G)=0$, a contradiction. $\Box$
Can we expect a lower sufficient minimum degree if we drop the condition on the size of the $(k+1)$-connected subgraph? At least, it does not matter essentially whether we demand size $2k$ or only size $\frac{4}{3}k$:
Theorem 2
  1. (i)
    There exists a graph with minimum degree 3k33k-3 without a $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices.
  2. (ii)
    There exists a graph with minimum degree $\big{(}3-\frac{1}{6}\big{)}k-\mathcal{O}(1)$ without a $(k+1)$-connected subgraph.
Proof.  We present a construction for (i), which we will modify for (ii) at the end.
The construction works in two steps. First, we construct a graph $H$ without a $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices in which there is a set $X$ of 2k12k-1 vertices of degree at least 2k22k-2 and all vertices in $V(H)\setminus X$ have degree at least 3k33k-3.
Let $a,b,c\in\mathbb{N}$ which differ by at most $1$ with a+b+c=k1a+b+c=k-1. Let $V_{0}$ be an anticlique on $k$ vertices and let $V_{1},\dots,V_{l}$ be disjoint cliques disjoint to $V_{0}$ on k1k-1 vertices, where $l=3+2a+b$. We partition each $V_{j}$ for $j\geq 1$ into three cliques $V_{j}=A_{j}\cup B_{j}\cup C_{j}$ with $|A_{j}|=a$, $|B_{j}|=b$ and $|C_{j}|=c$.
Let $w_{1},\dots,w_{2a+b}$ be a list of the vertices in $W:=A_{1}\cup B_{1}\cup A_{2}$. For $j\in\{1,2,3\}$, we add all possible edges between $V_{j}$ and $W_{j}:=V_{0}$. For $j\geq 4$, we add all possible edges between $V_{j}$ and the $k$ vertices in  $W_{j}\leavevmode\nobreak\ :=\leavevmode\nobreak\ A_{j-1}\,\cup\,B_{j-2}\,\cup\,C_{j-3}\,\cup\,\{w_{j-3}\}$. For $j=l$, this is shown in Figure 1.
Figure 1:  The graph $H$. Only the edges from $V_{0}$ and $V_{l}$ are shown.
Note that there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$. Let $H$ be the graph obtained this way and let $X:=C_{l-2}\cup B_{l-1}\cup C_{l-1}\cup V_{l}$. We have that
$|X|\leavevmode\nobreak\ =\leavevmode\nobreak\ c+b+c+(k-1)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 1+a+b+c+(k-1)\leavevmode\nobreak\ =\leavevmode\nobreak\ 2k-1\,.$
We first check the claim on the degrees: All vertices in $V_{0}$ have the neighborhood $V_{1}\cup V_{2}\cup V_{3}$ of size 3(k1)3(k-1). The vertices in some $V_{j}$ for $j\geq 1$ have edges to the other k2k-2 vertices in $V_{j}$ and to the $k$ vertices in $W_{j}$. Hence, the vertices in $X$ have degree 2k22k-2. Additionally, all vertices in $(V_{1}\cup\dots\cup V_{l})\setminus X$ are contained in some set $W_{j}$, so they have k1k-1 additional neighbors in $V_{j}$, which adds up to a degree of 3k33k-3.
We now prove inductively for $j\in\{3,4,\dots,l\}$ that there is no $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices in $G_{j}:=H[V_{0}\cup\dots\cup V_{j}]$.
For $j=3$ this is true, since $G_{3}$ is separated by removing the $k$ vertices from $V_{0}$, so the vertex set of a $(k+1)$-connected subgraph must be completely contained in $V_{0}\cup V_{1}$, $V_{0}\cup V_{2}$ or $V_{0}\cup V_{3}$. But in each of these cases, the vertices in $V_{0}$ have degree at most $k$ and $V_{1}$, $V_{2}$ or $V_{3}$ by themselves are too small for containing a $(k+1)$-connected subgraph.
Now let $j\geq 4$ and assume that there is a $(k+1)$-connected subgraph $G^{\prime}$ on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices in $G_{j}$. Since the neighborhood $W_{j}$ of $V_{j}$ in $G_{j}$ consists of only $k$ vertices, $G^{\prime}$ is either a subgraph of $G_{j}-V_{j}=G_{j-1}$ or of $G_{j}[V_{j}\cup W_{j}]$. The first case is prevented by the induction hypothesis, so assume the second. Since there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$, and those three are separated after deleting the set $V_{j}\cup\{w_{j-3}\}$ of $k$ vertices, $V(G^{\prime})$ must be completely contained in one of the sets $V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j-3}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j-3}\}\cup C_{j-3}$. But all of them have size at most $(k-1)+1+\big{\lceil}\frac{k-1}{3}\big{\rceil}\leq\big{\lceil}\frac{4}{3}k\big{\rceil}$, a contradiction.
This proves the claimed properties of $H$. We now inductively construct a series $H_{0},H_{1},\dots$ of graphs without a $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices such that in $H_{j}$ there is a set $X_{j}$ of $(2k-2^{j})\lor 0$ vertices of degree 2k22k-2 and all vertices in $V(H_{j})\setminus X_{j}$ have degree at least 3k33k-3. Then for $j\geq\log_{2}2k$, $H_{j}$ has minimum degree at least 3k33k-3.
Let $H_{0}=H$ and $X_{0}=X$ (possibly replenished to size 2k12k-1), which satisfy the conditions. Now assume we have already constructed $H_{j},X_{j}$ and we want to construct $H_{j+1},X_{j+1}$. Let $Y_{j}\subseteq X_{j}$ with $|Y_{j}|=k\land|X_{j}|$. Let $H^{\prime}_{j}$ be a copy of $H_{j}$ with $V(H_{j})\cap V(H^{\prime}_{j})=Y_{j}$ and $H_{j}[Y_{j}]=H^{\prime}_{j}[Y_{j}]$. We define $H_{j+1}=\big{(}V(H_{j})\cup V(H^{\prime}_{j}),\,E(H_{j})\cup E(H^{\prime}_{j})\big{)}$. Since the subgraphs $H_{j}$ and $H^{\prime}_{j}$ of $H_{j+1}$ intersect only in the vertex set $Y_{j}$ of at most $k$ vertices and contain no $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices, neither $H_{j+1}$ does. For every $y\in Y_{j}$, we have
$d_{H_{j+1}}(y)\leavevmode\nobreak\ =\leavevmode\nobreak\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2(k-1)+2(k-1)-(k-1)\leavevmode\nobreak\ =\leavevmode\nobreak\ 3(k-1)\,.$
Also the vertices that are not contained in $X_{j}$ or its copy, have degree at least 3k33k-3. The remaining vertices from $X_{j}\setminus Y_{j}$ and its copy form the new set $X_{j+1}$ of size
$|X_{j+1}|\leavevmode\nobreak$ = $=\leavevmode\nobreak 2|X_{j}\setminus Y_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak 2(|X_{j}|-k\land|X_{j}|)\leavevmode\nobreak\ =\leavevmode\nobreak 2(|X_{j}|-k)\lor 0$
= $=\leavevmode\nobreak 2\big{(}(2k-2^{j})\lor 0-k\big{)}\lor 0\leavevmode\nobreak\ =\leavevmode\nobreak 2(2k-2^{j}-k)\lor 0\leavevmode\nobreak\ =\leavevmode\nobreak\ (2k-2^{j+1})\lor 0\,.$
This finishes the proof of (i).
For (ii), we replace the complete graphs on each of the sets $A_{j},B_{j},C_{j}$ for $j\geq 1$ by the disjoint union of two complete graphs of size as equal as possible. By doing so, we loose in $H_{0}=H$ a degree $\frac{1}{6}k+\mathcal{O}(1)$ at every vertex. For the construction of $H_{1}$, we may assume $c\leq a$, and we let $Y_{0}=V_{l}\subseteq X$. Hence, $|X_{1}|$ consists of $C_{l-2}\cup B_{l-1}\cup C_{l-1}$ and its copy, which are $2(2c+b)\leq 2(a+b+c)=2k-2$ vertices. Doing so guarantees $d_{H_{j}[Y_{j}]}(y)\leq k-\frac{1}{6}k+\mathcal{O}(1)$ for every $j\geq 0$, which is needed to show for $y\in Y_{j}$ that
$d_{H_{j+1}}(y)\leavevmode\nobreak$ $=\leavevmode\nobreak\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)$
$\geq\leavevmode\nobreak 2\big{(}2k-\tfrac{1}{6}k-\mathcal{O}(1)\big{)}-\big{(}k-\tfrac{1}{6}k+\mathcal{O}(1)\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{(}3-\tfrac{1}{6}\big{)}k-\mathcal{O}(1)\,.$
The proof of the non-existence of a $(k+1)$-connected subgraph first works as for (i): The vertex set of a $(k+1)$-connected subgraph must be completely contained in one of the sets $V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j-3}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j-3}\}\cup C_{j-3}$, for some $j\geq 4$. Assume $V(G^{\prime})\subseteq V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$. The two components of $G[A_{j-1}]$ can be separated by deleting the other $k$ vertices. Hence $G^{\prime}$ cannot contain vertices from both components. The same holds for $A_{j}\subseteq V_{j}$. Together, $V(G^{\prime})\leq 2\big{\lceil}\frac{a}{2}\big{\rceil}+b+c+1\leq 1+a+b+c+1=k+1$, which is too low for being $(k+1)$-connected. The other two cases work similarly. $\Box$
Conjecture
There is a $C<3$ such that, for all $k\in\mathbb{N}$, every graph of minimum degree at least $Ck$ has a $(k+1)$-connected subgraph.
By Theorem On highly (vertex-)connected subgraphs, a suitable constant must satisfy $C\geq 3-\frac{1}{6}$. The conjecture is posed in a way that allows to assume that $k>k_{0}$ for some sufficiently large $k_{0}$: For $k\leq k_{0}$, we have $\big{(}3-\frac{1}{k_{0}}\big{)}k\,\geq\,\big{(}3-\frac{1}{k}\big{)}k\,=\,3k-1$, so Theorem On highly (vertex-)connected subgraphs suffices.
As another modification of Mader’s Conjecture, we consider triangle-free graphs.
Theorem 3
Every triangle-free graph with average degree at least $2k$ has a $(k+1)$-connected subgraph (on at least $2(k+1)$ vertices).
Proof.  First, we prove the well-known statement that, if $G$ is a triangle-free graph on $2l$ vertices, then $e(G)\leq l^{2}$. This is true for $l=1$. If $l>1$ and $E(G)\neq\emptyset$, consider an edge $vw\in E(G)$. Since $G$ is triangle-free, the neighborhoods of $v$ and $w$ are disjoint and hence $d_{G}(v)+d_{G}(w)\leq v(G)=2l$. This yields $e(G)=e(G-v-w)+d_{G}(v)+d_{G}(w)-1\leq(l-1)^{2}+2l-1=l^{2}$.
Now, let $G$ be a triangle-free graph with $v(G)\geq 2k$ without a $(k+1)$-connected subgraph. We prove by induction on $v(G)$ that $e(G)\leq k\big{(}v(G)-k\big{)}$. By the consideration above, the claim holds for $v(G)=2k$.
Now assume $v(G)>2k$. $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$ (see Theorem On highly (vertex-)connected subgraphs). By symmetry, we may assume that $v(A)\geq v(B)$.
Case I:  $v(B)\leq 2k$.
Assume that all vertices in $V(B)\setminus V(A)$ have degree greater than $k$. Let $b\in V(B)\setminus V(A)$. Since $|V(A)\cap V(B)|=k$, $b$ has a neighbor $c\in V(B)\setminus V(A)$. Since $B$ is triangle-free, the neighborhoods of $b$ and $c$ are disjoint subsets of $V(B)$. By the assumption, both contain at least $k+1$ vertices, which contradicts $v(B)\leq 2k$.
Hence, there is a vertex $b\in V(B)\setminus V(A)$ of degree at most $k$. By using the induction hypothesis for GbG-b,
$e(G)\leavevmode\nobreak\ =\leavevmode\nobreak\ e(G-b)+d_{G}(b)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ e(G-b)+k\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\big{(}v(G)-1-k\big{)}+k\leavevmode\nobreak\ =\leavevmode\nobreak\ k\big{(}v(G)-k\big{)}\,.$
Case II:  $v(B)>2k$.
Then also $v(A)>2k$, so we may use the induction hypothesis for both $A$ and $B$. Hence,
$e(G)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ e(A)+e(B)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\big{(}v(A)-k\big{)}+k\big{(}v(B)-k\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ k\big{(}v(G)-k\big{)}\,.$
This proves the claim.
Hence, if $G$ is a triangle-free graph with average degree at least $2k$, then $v(G)\geq 2k$ and $e(G)\geq kv(G)>k\big{(}v(G)-k\big{)}$. Hence, $G$ contains a $(k+1)$-connected subgraph $G^{\prime}$. Since $G^{\prime}$ is also triangle-free, for some edge $vw\in E(G^{\prime})$, the neighborhoods of $v$ and $w$ are disjoint. This shows $v(G^{\prime})\geq d_{G^{\prime}}(v)+d_{G^{\prime}}(w)\geq 2(k+1)$. $\Box$
This Theorem is best possible, even in the sense of minimum degree condition: There exists a bipartite graph with minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. One can build such a graph by gluing together copies of the bipartite graph $H_{0}:=K_{k,2k-1}$ similarly to the final step of the proof of Theorem On highly (vertex-)connected subgraphs.
On highly edge-connected subgraphs
Mader proved in [3] that every simple graph $G$ with $e(G)>k\big{(}v(G)-\frac{k+1}{2}\big{)}$ and $v(G)\geq k$ contains a $(k+1)$-edge-connected subgraph. This condition is sharp for every $k$ and every $v(G)$: Consider the graph $G$ on a clique $A$ on $k$ vertices and an anticlique on $V(G)\setminus A$ with all possible edges between $A$ and $V(G)\setminus A$. Clearly, $G$ has not even a subgraph of minimum degree at least $k+1$. Note that, when $v(G)$ tends to $\infty$, the average degree of $G$ tends to $2k$.
We first show that an average degree of at least $2k$ has some stronger implications. For simple graphs in particular, it implies the existence of a $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
Theorem 4
Let $G$ be a (not necessarily simple) graph with e(G)>k(v(G)1)e(G)\,>\,k\,(v(G)-1). Then $G$ contains a subgraph $G^{\prime}$ with
  1. (i)
    $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$.
  2. (ii)
    $G^{\prime}$ is $(k+1)$-edge-connected.
  3. (iii)
    $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees.
  4. (iv)
    If $G$ is simple, then $v(G^{\prime})>2k$.
Proof.  Let $G^{\prime}$ be a smallest nonempty induced subgraph of $G$ with $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$. Let $\mathcal{P}$ be a partition of $V(G^{\prime})$ in more than one class. Then for all $A\in\mathcal{P}$, $e(G[A])\leq k(|A|-1)$. Hence
$e(\mathcal{P})\leavevmode\nobreak$ $=\leavevmode\nobreak\ e(G^{\prime})\,-\sum_{A\in\mathcal{P}}e(G[A])\leavevmode\nobreak\ >\leavevmode\nobreak\ k(v(G^{\prime})-1)\,-\sum_{A\in\mathcal{P}}k(|A|-1)$
$=\leavevmode\nobreak\ k\Big{(}v(G)-1+|\mathcal{P}|-\sum_{A\in\mathcal{P}}|A|\Big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ k(|\mathcal{P}|-1)\,.$
In particular, it follows from the case $|\mathcal{P}|=2$, that $G^{\prime}$ is $(k+1)$-edge-connected. Moreover, by the Theorem of Tutte [4] / Nash-Williams [5], $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees. If $G$ and hence $G^{\prime}$ is simple, then $G^{\prime}$ has more than $2k$ vertices, since otherwise $e(G^{\prime})\,\leq\,\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)\,\leq\,k\,(v(G^{\prime})-1)$. $\Box$
In particular, Theorem On highly edge-connected subgraphs applies to graphs of minimum degree at least $2k$. Even the minimum degree version is best possible in some senses:
  1. (1)
    The complete graph on $2k$ vertices has minimum degree 2k12k-1 but no $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
  2. (2)
    Every (2k1)(2k-1)-regular simple graph $G$ without a complete subgraph on $2k$ vertices does not contain a subgraph $G^{\prime}$ with $k$ pairwise edge-disjoint spanning trees, because every subgraph $G^{\prime}$ has less than the necessary $k(v(G^{\prime})-1)$ edges:
    • If $v(G^{\prime})<2k$, then $e(G^{\prime})\leq\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)<k(v(G^{\prime})-1)$.
    • If $v(G^{\prime})=2k$, then $e(G^{\prime})<\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)=k(v(G^{\prime})-1)$.
    • If $v(G^{\prime})>2k$, then $e(G^{\prime})\leq\frac{1}{2}(2k-1)v(G^{\prime})=kv(G^{\prime})-\frac{1}{2}v(G^{\prime})<k(v(G^{\prime})-1)$.
  3. (3)
    There exists a multi-graph with minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph: Consider the graph $G$ on vertex set $\{1,\dots,2(k+1)\}$ with k1k-1 parallel edges between each two consecutive vertices $i$ and $i+1$, one additional edge between $1$ and each of the $k$ vertices $\{2,\dots,k+1\}$, and between $2(k+1)$ and each of the $k$ vertices $\{k+2,\dots,2k+1\}$. Clearly, all vertices have degree 2k12k-1. Between the two vertex sets $\{1,\dots,k+1\}$ and $\{k+2,\dots,2k+2\}$ there are only k1k-1 edges. Hence, if $G^{\prime}$ is a $(k+1)$-connected subgraph of $G$, then either $V(G^{\prime})\subseteq\{1,\dots,k+1\}$ or $V(G^{\prime})\subseteq\{k+2,\dots,2k+2\}$. In both cases, there is a vertex of degree at most $k$, $\max V(G^{\prime})$ or $\min V(G^{\prime})$, respectively. This is a contradiction.
  4. (4)
    A similar construction can be used for simple graphs. Together with ideas from Theorem On highly edge-connected subgraphs, one can show that for $k\leq 5$, a minimum degree 2k12k-1 is still best possible: One can build graphs of minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. Because the construction allows no generalization, we won’t show it here. Instead, we now consider large $k$.
We now prove that, for simple graphs and for $k\to\infty$, already a minimum degree $\big{(}1+o(1)\big{)}k$ is sufficient for a $(k+1)$-edge-connected subgraph.
Let $\alpha\in\big{(}0,\frac{1}{2}\big{)}$ be fixed. All of the following definitions will depend on $k$, which will not be noticeable in the notations. Let $m=m(k)\in[0,k]$ not necessarily an integer, which we will choose later.
Let $G$ be a graph and $v\in V(G)$ of degree $d_{G}(v)$. We define a penalty
$m_{G}(v)\leavevmode\nobreak\ :=\leavevmode\nobreak\ 0\lor\big{(}k+m-d_{G}(v)\big{)}\land m\leavevmode\nobreak\ \in[0,m]\,.$
We consider the following weighted penalty of the graph $G$:
[…]
The sum of the first $l$ weights can be upper-bounded as follows:
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha j^{\alpha-1}dx\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ \int_{0}^{l}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ l^{\alpha}\,.$
Similarly, we obtain as a lower bound:
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha j^{\alpha-1}dx\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ \int_{1}^{l+1}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ \geq\leavevmode\nobreak\ l^{\alpha}-1\,.$
Lemma On highly edge-connected subgraphsa
Let $G$ be a graph that contains a vertex $w$ of degree at most $k$. Then
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-k^{\alpha}\,.$
Proof.  Let $v_{1},\dots,v_{v(G)-1}$ be a list of $V(G-w)=V(G)\setminus\{w\}$ that attains
$M(G-w)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G-w}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)-1\}$, we have $d_{G}(v_{j})-d_{G-w}(v_{j})\in\{0,1\}$ as $G$ is simple. Since the function $x\mapsto 0\lor\big{(}(k+m)-x\big{)}\land m$ is monotone decreasing and non-expanding, this yields $\Delta_{j}:=m_{G-w}(v_{j})-m_{G}(v_{j})\in[0,1]$ and
$\sum_{j=1}^{v(G)-1}\Delta_{j}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}d_{G}(v_{j})-d_{G-w}(v_{j})\leavevmode\nobreak\ =\leavevmode\nobreak\ d_{G}(w)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\,.$
Consider the list $v_{1},\dots,v_{v(G)-1},w$ of the vertices of $G$. This yields
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G}(v_{v(G)-1})+\alpha v(G)^{\alpha-1}m_{G}(w)$
$>\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\big{(}m_{G-w}(v_{j})-\Delta_{j}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ M(G-w)-\sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)-1}\Delta_{j}\leq k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, the weighted sum over the $\Delta_{j}$ is maximized if $\Delta_{j}=1$ for $j\leq k$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-\sum_{j=1}^{k}\alpha j^{\alpha-1}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-k^{\alpha}\,$,
using the upper bound of the first $k$ weights. $\Box$
Lemma On highly edge-connected subgraphsb
Let $G$ be a graph that contains a set $S$ of at most $k$ edges such that GSG-S is not connected, so it decomposes into non-empty subgraphs $A$ and $B$, that is, $V(G)$ is partitioned into $V(A)\cup V(B)$ and $E(G)\setminus S$ is partitioned into $E(A)\cup E(B)$. Then
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\,.$
Proof.  Let $a_{1},\dots,a_{v(A)}$ and $b_{1},\dots,b_{v(B)}$ be lists of $V(A)$ and $V(B)$ that attain
$M(A)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha j^{\alpha-1}m_{A}(a_{j})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ M(B)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(B)}\alpha j^{\alpha-1}m_{B}(b_{j})\,.$
Clearly $d_{A}(a_{j})=d_{G-S}(a_{j})$ and hence $m_{A}(a_{j})=m_{G-S}(a_{j})$. Similarly $m_{B}(b_{j})=m_{G-S}(b_{j})$. We may assume by symmetry that $v(A)\leq v(B)$. Consider the following list of V(GS)=V(G)V(G-S)=V(G):
$a_{1},b_{1},a_{2},b_{2},\dots,a_{v(A)},b_{v(A)},\dots,b_{v(B)}\,.$
We obtain
$M(G-S)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha(2j-1)^{\alpha-1}m_{A}(a_{j})+\sum_{j=1}^{v(A)}\alpha(2j)^{\alpha-1}m_{B}(b_{j})+\hskip-10.00002pt\sum_{j=v(A)+1}^{v(B)}\hskip-10.00002pt\alpha(v(A)+j)^{\alpha-1}m_{B}(b_{j})$
$\geq\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha(2j)^{\alpha-1}m_{A}(a_{j})+\sum_{j=1}^{v(B)}\alpha(2j)^{\alpha-1}m_{B}(b_{j})\leavevmode\nobreak\ =\leavevmode\nobreak 2^{\alpha-1}M(A)+2^{\alpha-1}M(B)\,.$
Let $v_{1},\dots,v_{v(G)}$ be a list of V(GS)=V(G)V(G-S)=V(G) that attains
$M(G-S)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)\}$, we have $d_{G}(v_{j})\geq d_{G-S}(v_{j})$, and hence $\Delta_{j}:=m_{G-S}(v_{j})-m_{G}(v_{j})\in[0,m]$. It is
$\sum_{j=1}^{v(G)}\Delta_{j}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{v(G)}d_{G}(v_{j})-d_{G-S}(v_{j})\leavevmode\nobreak\ =\leavevmode\nobreak\ 2|S|\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2k\,.$
By considering the same list as a list of $V(G)$, we obtain
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\big{(}m_{G-S}(v_{j})-\Delta_{j}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ M(G-S)-\sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)}\Delta_{j}\leq 2k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, we obtain an upper bound of the weighted sum over the $\Delta_{j}$ with $\Delta_{j}=m$ for $j\leq\big{\lceil}\frac{2k}{m}\big{\rceil}$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-S)-\sum_{j=1}^{\big{\lceil}\frac{2k}{m}\big{\rceil}}\alpha j^{\alpha-1}m\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\,$,
using the upper bound of the first $\big{\lceil}\frac{2k}{m}\big{\rceil}\leq\frac{4k}{m}$ weights. $\Box$
Theorem 5
For every $\alpha>0$, there is some $k_{0}$ such that, for all $k\geq k_{0}$, every graph with minimum degree at least $k+k^{\frac{1}{2}+\alpha}$ has a $(k+1)$-edge-connected subgraph.
Proof.  Let $\alpha\in\big{(}0,\frac{1}{2}\big{)}$$m=k^{\frac{1}{2}+\alpha}\leq k$  and  $\mu=\frac{1}{2}m\sqrt{k}^{\alpha}$.
Let $G$ be a graph with $v(G)>k$ without a $(k+1)$-edge-connected subgraph. We prove by induction on $v(G)$ that $M(G)\geq\mu$.
A sequence of vertices $w_{1},\dots,w_{l}\in V(G)$ is called detachable from $G$, if for each $j\in\{1,\dots,l\}$, $w_{j}$ has degree at most $k$ in $G-w_{1}-\dots-w_{j-1}$. We set
$z(G)\leavevmode\nobreak\ :=\leavevmode\nobreak\ \max\big{\{}l\leavevmode\nobreak\ :\leavevmode\nobreak\ w_{1},\dots,w_{l}\in V(G)\text{ is detachable from }G\big{\}}\,.$
Case I:  $z(G)\geq\sqrt{k}$.
Then, there is a detachable sequence $w_{1},\dots,w_{\big{\lceil}\sqrt{k}\big{\rceil}}\in V(G)$. For each $j\in\{1,\dots,l\}$$d_{G}(w_{j})\leq k+(j-1)$ and hence $m_{G}(w_{j})\geq m-(j-1)\geq m-(\big{\lceil}\sqrt{k}\big{\rceil}-1)\geq m-\sqrt{k}\geq\frac{2}{3}m$,for $k$ large enough. This implies, again for $k$ large enough,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{\big{\lceil}\sqrt{k}\big{\rceil}}\alpha j^{\alpha-1}m_{G}(v_{j})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{2}{3}m\sum_{j=1}^{\big{\lceil}\sqrt{k}\big{\rceil}}\alpha j^{\alpha-1}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{2}{3}m(\sqrt{k}^{\alpha}-1)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}m\sqrt{k}^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mu\,.$
Case II:  $z(G)<\sqrt{k}$.
Then, there is a longest detachable sequence $w_{1},\dots,w_{z(G)}\in V(G)$. Hence, in the graph $G^{\prime}=G-w_{1}-\dots-w_{z(G)}$, all vertices have degree >k>k. Since $G^{\prime}$ is a non-empty subgraph of $G$, it is not $(k+1)$-edge-connected. Hence, there is a set $S$ of at most $k$ edges such that $G^{\prime}-S$ decomposes into two non-empty subgraphs $A$ and $B$. We may assume that all edges of $S$ are between $V(A)$ and $V(B)$.
Since $d_{G^{\prime}}(a)>k$ for all $a\in V(A)$, we have
$(v(A)-1)k\leavevmode\nobreak\ =\leavevmode\nobreak\ v(A)k-k\leavevmode\nobreak\ <\sum_{a\in V(A)}d_{G^{\prime}}(a)\leavevmode\nobreak\ -|S|\leavevmode\nobreak\ =\sum_{a\in V(A)}d_{A}(a)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ v(A)(v(A)-1)\,$,
which yields $v(A)>k$, and similarly $v(B)>k$. Like $G$, also its subgraphs $A$ and $B$ do not contain a $(k+1)$-edge-connected subgraph, so we can apply the induction hypothesis to both $A$ and $B$. Together with Lemma On highly edge-connected subgraphsb, we obtain
$M(G^{\prime})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-m\,o(\sqrt{k})^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-o(\mu)\,.$
Finally, applying Lemma On highly edge-connected subgraphsa on each of the vertices $w_{1},\dots,w_{z(G)}$ yields
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w_{1}-\dots-w_{z(G)})-z(G)k^{\alpha}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G^{\prime})-k^{\frac{1}{2}}k^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-o(\mu)\,.$
Hence, for $k$ large enough, $M(G)\geq\mu$.
This finishes the proof: If $G$ has minimum degree at least $k+m$, then $v(G)>k$ and $M(G)=0$, a contradiction. $\Box$
The condition $\alpha>0$ is sharp:
Theorem 6
For every constant $C$, there exists a $k\in\mathbb{N}$ and a graph with minimum degree at least $k+C\sqrt{k}$ without a $(k+1)$-edge-connected subgraph.
Proof.  We may assume that $C\in\mathbb{N}$ and we choose $k$ as a square such that $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$. Further let $k$ be large enough such that $2^{2C^{2}}C\leq\sqrt{k}$.
We prove for $j=0,\dots,2C^{2}$ that there is a graph $G_{j}$ and $A_{j}\subseteq V(G_{j})$ with
  1. (i)
    $G_{j}$ does not contain a $(k+1)$-edge-connected subgraph.
  2. (ii)
    $|A_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{j}(C\sqrt{k}+1)$.
  3. (iii)
    The vertices in $A_{j}$ have degree at least $k+\frac{1}{2C}j\sqrt{k}$.
  4. (iv)
    The vertices in $V(G_{j})\setminus A_{j}$ have degree at least $k+C\sqrt{k}$.
The claim follows with $j=2C^{2}$.
Define $G_{0}$ as a graph on $k+C\sqrt{k}+1$ vertices that contains an anticlique $A_{0}$ of $C\sqrt{k}+1$ vertices, a clique $V(G_{0})\setminus A_{0}$ on $k$ vertices and all possible edges between $A_{0}$ and $V(G_{0})\setminus A_{0}$. The vertices in $A_{0}$ have degree $k$ and those in $V(G_{0})\setminus A_{0}$ have degree $k+C\sqrt{k}$. A $(k+1)$-connected subgraph of $G_{0}$ cannot contain any vertex from $A_{0}$, but also $G_{0}-A_{0}$ has to few vertices to contain a $(k+1)$-connected subgraph.
Now let $0\leq j<2C^{2}$. Assume we have already constructed $G_{j},A_{j}$ and we want to construct $G_{j+1},A_{j+1}$. Note that
$|A_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{j}(C\sqrt{k}+1)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2^{j+1}C\sqrt{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2^{2C^{2}}C\sqrt{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\,.$
We build a graph $H_{j}$ from $G_{j}$ by adding an anticlique $B_{j}$ on $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$ vertices. From each vertex in $B_{j}$, we add one edge to every vertex in $A_{j}$ and $k-|A_{j}|$ more edges to other vertices from $G_{j}$. In $H_{j}$, the vertices in $B_{j}$ have degree $k$, the vertices in $A_{j}$ now have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ and all other vertices still have degree at least $k+C\sqrt{k}$. A $(k+1)$-connected subgraph of $H_{j}$ cannot contain any vertex from $B_{j}$, but also $G_{j}=H_{j}-B_{j}$ does not contain one by the assumption.
We build $G_{j+1}$ from $H_{j}$ and one disjoint copy $H^{\prime}_{j}$. Let $A^{\prime}_{j}$ and $B^{\prime}_{j}$ be the copies of $A_{j}$ and $B_{j}$ in $H^{\prime}_{j}$. From every vertex in $B_{j}$, we add $C\sqrt{k}$ edges to vertices in $H^{\prime}_{j}$. Similarly, from every vertex in $B^{\prime}_{j}$, we add edges to vertices in $H_{j}$ until there are $C\sqrt{k}$ many. All together, we have added at most $2\frac{1}{2C}\sqrt{k}C\sqrt{k}=k$ edges between $H_{j}$ and $H^{\prime}_{j}$. Hence, a $(k+1)$-edge-connected subgraph of $G_{j+1}$ must be completely contained in either $H_{j}$ or $H^{\prime}_{j}$, which we prevented with the construction.
We set $A_{j+1}:=A_{j}\cup A^{\prime}_{j}$ of size $2|A_{j}|=2^{j+1}(C\sqrt{k}+1)$. The vertices in $A_{j+1}$ still have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ as in $H_{j}$. All other vertices, including those in $B_{j}\cup B^{\prime}_{j}$, have degree at least $k+C\sqrt{k}$. This finishes the construction. $\Box$
For the proof, we need $C=\Theta(\sqrt{\log k})$, which might be sharp: For some $D$, a minimum degree of at least $k+D\sqrt{k\log k}$ might be sufficient for a $(k+1)$-edge-connected subgraph.
Application on graph decomposition
The Theorems from the present paper have some applications for other issues. We want to mention decomposition problems of the following type, which were first studied by Thomassen [6]: What is the lowest number $k=k(s,t)$ such that every kk-(edge-)connected graph $G$ admits a partition of $V(G)$ into $A$ and $B$ such that $G[A]$ is ss-(edge-)connected and $G[B]$ is tt-(edge-)connected?
Corollary.   Let $s,t\geq 1$.
  1. (i)
    Let $G$ be an $(s+t+1)$-connected graph of minimum degree at least 3s+3t13s+3t-1. Then $V(G)$ has a partition into two sets $A$ of size >2s>2s and $B$ of size >2t>2t such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected.
  2. (ii)
    Let $G$ be a triangle-free and $(s+t+1)$-connected graph of minimum degree at least $2s+2t$. Then $V(G)$ has a partition into two sets $A$ and $B$ such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected.
  3. (iii)
    There is a function $f(r)=\big{(}1+o(1)\big{)}r$ such that the following holds: Let $G$ be an $(s+t+1)$-edge-connected graph of minimum degree at least $f(s+t)$. Then $V(G)$ has a partition into two sets $A$ and $B$ such that $G[A]$ is $(s+1)$-edge-connected and $G[B]$ is $(t+1)$-edge-connected.
Proof.  The approach is due to Thomassen [6].
  • (i)
    $G$ has minimum degree at least (3s1)+(3t1)+1(3s-1)+(3t-1)+1. By a result of Stiebitz [7], $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ has minimum degree at least 3s13s-1 and $G[B^{\prime}]$ has minimum degree at least 3t13t-1. By Theorem On highly (vertex-)connected subgraphs, $G[A^{\prime}]$ has an $(s+1)$-connected subgraph on some vertex set $A$ of size >2s>2s and $G[B^{\prime}]$ has a $(t+1)$- connected subgraph on some vertex set $B$ of size >2t>2t. We now ignore $A^{\prime}$ and $B^{\prime}$.
    Among all disjoint $A,B\subseteq V(G)$ with $|A|>2s$ and $|B|>2t$ such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected, we choose $A$ and $B$ with minimum size of $X=V(G)\setminus(A\cup B)$. Assume $X\neq\emptyset$.
    Then $G[A\cup X]$ is not $(s+1)$-connected, that is, there is some $S\subseteq A\cup X$ of at most $s$ vertices such that $G[(A\cup X)\setminus S]$ decomposes into more than one connected component. One of them has to be disjoint from $A$. Let $Y\subseteq X$ be its vertex set.
    Now $G[B\cup Y]$ is not $(t+1)$-connected, that is, there is some $T\subseteq B\cup Y$ of at most $t$ vertices such that $G[(B\cup Y)\setminus T]$ decomposes into more than one connected component. One of them has to be disjoint from $B$. Let $Z\subseteq Y$ be its vertex set. The neighborhood of $Z$ is contained in $S\cup T$ of size at most $s+t$, a contradiction to the $(s+t+1)$-connectedness of $G$.
  • (ii)
    By a result of Kaneko [8] on triangle-free graphs, $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ has minimum degree at least $2s$ and $G[B^{\prime}]$ has minimum degree at least $2t$. Now use Theorem On highly (vertex-)connected subgraphs to obtain disjoint $(s+1)$- and $(t+1)$-connected subgraphs. Then proceed as before.
  • (iii)
    By Theorem On highly edge-connected subgraphs and Theorem On highly edge-connected subgraphs for small $k$, there exists a function $g(k)=\big{(}1+o(1)\big{)}k$,$g(k)<2k$ such that for every $k$, every graph of minimum degree at least $g(k)$ contains a $(k+1)$-edge-connected subgraph. We choose $f(r)=\max\big{\{}g(r-s)+g(s)+1\,|\,1\leq s\leq\frac{r}{2}\big{\}}$. We have that $f(r)=\big{(}1+o(1)\big{)}r$ for $r\to\infty$:
    If $s\geq\sqrt{r}$, then $g(r-s)+g(s)+1=\big{(}1+o(1)\big{)}(r-s)+\big{(}1+o(1)\big{)}s+1=\big{(}1+o(1)\big{)}r$.Otherwise, $g(r-s)+g(s)+1<\big{(}1+o(1)\big{)}(r-s)+2s+1\leq\big{(}1+o(1)\big{)}r+o(r)=\big{(}1+o(1)\big{)}r$.
    Now let $G$ be an $(s+t+1)$-edge-connected graph of minimum degree at least $f(s+t)\geq g(s)+g(t)+1$. Again by Stiebitz [7], $V(G)$ has a partition into two sets $A^{\prime}$ and $B^{\prime}$ such that $G[A^{\prime}]$ and $G[B^{\prime}]$ have minimum degree at least $g(s)$ and $g(t)$, respectively. We obtain disjoint $(s+1)$- and $(t+1)$-edge-connected subgraphs. Since Thomassen’s argument also works for edge-connectedness, we may proceed as before. $\Box$