### v2

# arXiv:2508.07997v2[math.CO] 02 Mar 2026

Abstract
Mader (1979) conjectured that an average degree of at least 3k13k-1 in a graph is sufficient for the existence of a $(k+1)$-connected subgraph. The following minimum degree version holds: Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices. Moreover, for triangle-free graphs, already an average degree of at least $2k$ is sufficient for a $(k+1)$-connected subgraph, which has at least $2(k+1)$ vertices.
For edge-connectedness (in simple graphs), we prove the following: Every graph of average degree at least $2k$ has a $(k+1)$-edge-connected subgraph on more than $2k$ vertices. Moreover, for every small $\alpha>0$ and for $k$ large enough in terms of $\alpha$, already a minimum degree of at least $k+k^{\frac{1}{2}+\alpha}=\big(1+o(1)\big)k$ is sufficient for a $(k+1)$-edge-connected subgraph.
It is shown that all of these results are sharp in some sense. The results may be applied to decompose graphs into two highly connected or edge-connected parts.
Sufficient minimum degree conditions for the existence
of highly connected or edge-connected subgraphs
Maximilian Krone
Technische Universität Ilmenau
On highly connected subgraphs
Mader conjectured in [1] that every graph with average degree 3k13k-1 has a $(k+1)$-connected subgraph. This conjecture turned out to be quite hard, since the best known sufficient linear factor was improved in many iterations, but never reached $3$. We consider another version of the problem by demanding even a minimum degree of 3k13k-1. In exchange for the stronger requirement, the $(k+1)$-connected subgraph is enforced to have a non-trivial size of more than $2k$ vertices. This bonus outcome cannot be expected from an average degree of 3k13k-1, since by a construction of Carmesin in [2], there are graphs of average degree $\big(3+\frac{1}{3}\big)k-\mathcal{O}(1)$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
Theorem 1
Let $k\geq 1$ be fixed. Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices.
Proof. Let $G$ be a graph with $v(G)>2k$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
Let $v\in V(G)$ be a vertex of degree $d_{G}(v)$. We consider a delimited penalty
$m_{G}(v)\penalty 10000\ :=\penalty 10000\ \min\Big\{\max\big\{(3k-1)-d_{G}(v),\,0\big\},\,k\Big\}\penalty 10000\ \in\{0,\dots,k\}\,.$
We prove by induction on $v(G)>2k$ that the total penalty of $G$ satisfies
$M(G)\penalty 10000\ :=\penalty 10000\ \sum_{v\in V(G)}m_{G}(v)\penalty 10000\ \geq\penalty 10000\ 2k^{2}\,.$
Since $v(G)>2k$, $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$, that is, two induced but not spanning subgraphs $A$ and $B$ with $V(A)\cup V(B)=V(G)$,  $E(A)\cup E(B)=E(G)$ and $|V(A)\cap V(B)|=k$. By symmetry, we may assume that $v(A)\geq v(B)$.
Let $a\in V(A)\setminus V(B)$, $b\in V(B)\setminus V(A)$ and $s\in V(A)\cap V(B)$. We have $d_{G}(a)=d_{A}(a)$ and hence $m_{G}(a)=m_{A}(a)$. Similarly, $m_{G}(b)=m_{B}(b)$.
Case I: $v(A)\leq 2k$ and $v(B)\leq 2k$.
We have $d_{G}(s)\leq v(G)-1$, and hence $m_{G}(s)\geq\min\big\{(3k-1)-(v(G)-1),k\big\}=3k-v(G)$.It is $m_{G}(a)=m_{G}(b)=k$. Hence
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{s}\big(3k-v(G)\big)+\sum_{a}k+\sum_{b}k\penalty 10000\ =\penalty 10000\ k(3k-v(G))+(v(G)-k)k\penalty 10000\ =\penalty 10000\ 2k^{2}\,.$
Case II: $v(A)>2k$ and $v(B)\leq 2k$.
We use the induction hypothesis for $A$. Since the function $x\mapsto\min\big\{\max\{3k-1\,-x,\,0\},\,k\big\}$ is non-expanding, we have $m_{A}(s)-m_{G}(s)\leq\big|d_{A}(s)-d_{G}(s)\big|\leq|V(G)\setminus V(A)|=v(B)-k$. Again, $m_{G}(b)=k$. Hence
[…]
Case III: $v(A)>2k$ and $v(B)>2k$.
We use the induction hypothesis for $A$ and $B$. We have $m_{G}(s)\geq 0\geq m_{A}(s)+m_{B}(s)-2k$. Hence
[…]
$2.\displaystyle=\penalty 10000\ M(A)+M(B)-k(2k)\penalty 10000\ \geq\penalty 100002k^{2}+2k^{2}-2k^{2}\penalty 10000\ =\penalty 100002k^{2}\,$.
This finishes the proof: If $G$ has minimum degree at least 3k13k-1, then $v(G)>2k$ and $M(G)=0$, a contradiction. $\Box$
Can we expect a lower sufficient minimum degree if we drop the condition on the size of the $(k+1)$-connected subgraph? At least, it does not matter essentially whether we demand size $2k$ or only size $\frac{4}{3}k$:
Theorem 2
  1. (i)
    There exists a graph with minimum degree 3k33k-3 without a $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices.
  2. (ii)
    There exists a graph with minimum degree $\big(3-\frac{1}{6}\big)k-\mathcal{O}(1)$ without a $(k+1)$-connected subgraph.
Proof. We present a construction for (i), which we will modify for (ii) afterwards.
The construction works in two steps. First, we construct a graph $H$ without a $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices in which there is a set $X$ of 2k12k-1 vertices of degree at least 2k22k-2 and all vertices in $V(H)\setminus X$ have degree at least 3k33k-3.
Let $a,b,c\in\mathbb{N}$ which differ by at most $1$ with a+b+c=k1a+b+c=k-1. Let $V_{0}$ be an anticlique on $k$ vertices and let $V_{1},\dots,V_{l}$ be disjoint cliques on k1k-1 vertices that are also disjoint to $V_{0}$, where $l=3+2a+b$. We partition each $V_{j}$ for $j\geq 1$ into three cliques $V_{j}=A_{j}\cup B_{j}\cup C_{j}$ with $|A_{j}|=a$, $|B_{j}|=b$ and $|C_{j}|=c$.
Let $w_{4},w_{5},\dots,w_{l}$ be a list of the l3=2a+bl-3=2a+b vertices in $W:=A_{1}\cup B_{1}\cup A_{2}$. For $j\in\{1,\dots,l\}$, we add all possible edges between $V_{j}$ and some set $W_{j}\subseteq V_{0}\cup\dots\cup V_{j-1}$ of $k$ vertices. We set $W_{j}=V_{0}$ for $j\in\{1,2,3\}$ and $W_{j}=A_{j-1}\,\cup\,B_{j-2}\,\cup\,C_{j-3}\,\cup\,\{w_{j}\}$ otherwise. For $j\in\{1,2,3,l\}$, this is shown in Figure 1.
Note that there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$. Let $H$ be the graph obtained this way and let $X:=C_{l-2}\cup B_{l-1}\cup C_{l-1}\cup V_{l}$. We have that
$|X|\penalty 10000\ =\penalty 10000\ c+b+c+(k-1)\penalty 10000\ \leq\penalty 10000\ 1+a+b+c+(k-1)\penalty 10000\ =\penalty 10000\ 2k-1\,.$
Figure 1: The graph $H$. Only the edges between $V_{j}$ and $W_{j}$ for $j\in\{1,2,3,l\}$ are shown.
We first check the claim on the degrees: All vertices in $V_{0}$ have the neighborhood $V_{1}\cup V_{2}\cup V_{3}$ of size 3(k1)3(k-1). The vertices in some $V_{j}$ for $j\geq 1$ have edges to the other k2k-2 vertices in $V_{j}$ and to the $k$ vertices in $W_{j}$. Hence, the vertices in $X$ have degree 2k22k-2. Additionally, all vertices in $(V_{1}\cup\dots\cup V_{l})\setminus X$ are contained in some set $W_{j}$, so they have k1k-1 additional neighbors in $V_{j}$, which adds up to a degree of 3k33k-3.
We now prove inductively for $j\in\{3,4,\dots,l\}$ that there is no $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices in $G_{j}:=H[V_{0}\cup\dots\cup V_{j}]$.
For $j=3$ this is true, since $G_{3}$ is separated by removing the $k$ vertices from $V_{0}$, so the vertex set of a $(k+1)$-connected subgraph must be completely contained in $V_{0}\cup V_{1}$, $V_{0}\cup V_{2}$ or $V_{0}\cup V_{3}$. But in each of these cases, the vertices in $V_{0}$ have degree at most $k$ and $V_{1}$, $V_{2}$ or $V_{3}$ by themselves are too small for containing a $(k+1)$-connected subgraph.
Now let $j\geq 4$ and assume that there is a $(k+1)$-connected subgraph $G^{\prime}$ on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices in $G_{j}$. Since the neighborhood $W_{j}$ of $V_{j}$ in $G_{j}$ consists of only $k$ vertices, $G^{\prime}$ is either a subgraph of $G_{j}-V_{j}=G_{j-1}$ or of $G_{j}[V_{j}\cup W_{j}]$. The first case is prevented by the induction hypothesis, so assume the second. Since there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$, and those three are separated after deleting the set $V_{j}\cup\{w_{j}\}$ of $k$ vertices, $V(G^{\prime})$ must be completely contained in one of the sets $V_{j}\cup\{w_{j}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j}\}\cup C_{j-3}$. But all of them have size at most $(k-1)+1+\big\lceil\frac{k-1}{3}\big\rceil\leq\big\lceil\frac{4}{3}k\big\rceil$, a contradiction.
This proves the claimed properties of $H$. We now inductively construct a series $H_{0},H_{1},\dots$ of graphs without a $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices such that in $H_{j}$ there is a set $X_{j}$ of $\max\{2k-2^{j},0\}$ vertices of degree at least 2k22k-2 and all vertices in $V(H_{j})\setminus X_{j}$ have degree at least 3k33k-3. Then for $j\geq\log_{2}2k$, $H_{j}$ has minimum degree at least 3k33k-3.
Let $H_{0}=H$ and $X_{0}=X$ (possibly replenished to size 2k12k-1), which satisfy the conditions. Now assume we have already constructed $H_{j},X_{j}$ and we want to construct $H_{j+1},X_{j+1}$. Let $Y_{j}\subseteq X_{j}$ with $|Y_{j}|=\min\{k,|X_{j}|\}$. Let $H^{\prime}_{j}$ be a copy of $H_{j}$. We define $H_{j+1}$ as the union of $H_{j}$ and $H^{\prime}_{j}$ when we identify $Y_{j}$ and its copy in $H^{\prime}_{j}$, that is, the subgraphs $H_{j}$ and $H^{\prime}_{j}$ of $H_{j+1}$ intersect only in the vertex set $Y_{j}$. Since $|Y_{i}|\leq k$ and both $H_{j}$ and $H^{\prime}_{j}$ contain no $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices, neither $H_{j+1}$ does. For every $y\in Y_{j}$, we have
$d_{H_{j+1}}(y)\penalty 10000\ =\penalty 10000\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)\penalty 10000\ \geq\penalty 10000\ 2(k-1)+2(k-1)-(k-1)\penalty 10000\ =\penalty 10000\ 3(k-1)\,.$
Also the vertices that are not contained in $X_{j}$ or its copy in $H_{j}^{\prime}$ have degree at least 3k33k-3. The remaining vertices from $X_{j}\setminus Y_{j}$ and its copy in $H_{j}^{\prime}$ form the new set $X_{j+1}$ of size
[…]
This finishes the proof of (i).
For (ii), we replace the complete graphs on each of the sets $A_{j},B_{j},C_{j}$ for $j\geq 1$ by the disjoint union of two complete graphs of size as equal as possible. By doing so, we loose in $H_{0}=H$ a degree $\frac{1}{6}k+\mathcal{O}(1)$ at every vertex. One observes inductively for $j=0,1,\dots$ that $d_{H_{j}[Y_{j}]}(v)\leq d_{H_{j}[X_{j}]}(v)\leq k-\frac{1}{6}k+\mathcal{O}(1)$ for every vertex $v\in Y_{j}$. Hence indeed
[…]
The proof of the non-existence of a $(k+1)$-connected subgraph first works as for (i): The vertex set of a $(k+1)$-connected subgraph $G^{\prime}$ must be completely contained in one of the sets $V_{j}\cup\{w_{j}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j}\}\cup C_{j-3}$, for some $j\geq 4$. Assume $V(G^{\prime})\subseteq V_{j}\cup\{w_{j}\}\cup A_{j-1}$. The two components of $H[A_{j-1}]$ can be separated by deleting the set $V_{j}\cup\{w_{j}\}$ of $k$ vertices. Hence $G^{\prime}$ cannot contain vertices from both components. The same holds for $A_{j}\subseteq V_{j}$. Together, $V(G^{\prime})\leq 2\big\lceil\frac{a}{2}\big\rceil+b+c+1\leq 1+a+b+c+1=k+1$, which is too low for being $(k+1)$-connected. The other two cases work similarly. $\Box$
Conjecture
There is a $C<3$ such that, for all $k\in\mathbb{N}$, every graph of minimum degree at least Ck1Ck-1 has a $(k+1)$-connected subgraph.
By Theorem On highly connected subgraphs, a suitable constant must satisfy $C\geq 3-\frac{1}{6}$. The conjecture is posed in a way that allows to assume that $k\geq k_{0}$ for some sufficiently large $k_{0}$: For $k<k_{0}$, $\big\lceil\big(3-\frac{1}{k_{0}}\big)k-1\big\rceil\,=\,\big\lceil 3k-\frac{k}{k_{0}}-1\big\rceil\,=\,3k-1$, so Theorem On highly connected subgraphs suffices for $C=3-\frac{1}{k_{0}}<3$.
As another modification of Mader’s Conjecture, we consider triangle-free graphs.
Theorem 3
Let $k\geq 1$. Every triangle-free graph with average degree at least $2k$ has a $(k+1)$-connected subgraph (on at least $2(k+1)$ vertices).
Proof. First, we prove the well-known statement that, if $G$ is a triangle-free graph on $2l$ vertices, then $e(G)\leq l^{2}$. This is true for $l=1$. If $l>1$ and $E(G)\neq\emptyset$, consider an edge $vw\in E(G)$. Since $G$ is triangle-free, the neighborhoods of $v$ and $w$ are disjoint and hence $d_{G}(v)+d_{G}(w)\leq v(G)=2l$. This yields $e(G)=e(G-v-w)+d_{G}(v)+d_{G}(w)-1\linebreak\leq(l-1)^{2}+2l-1=l^{2}$.
Now, let $G$ be a triangle-free graph with $v(G)\geq 2k$ without a $(k+1)$-connected subgraph. We prove by induction on $v(G)$ that $e(G)\leq k\big(v(G)-k\big)$. By the consideration above, the claim holds for $v(G)=2k$.
Now assume $v(G)>2k$. $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$ (see Theorem On highly connected subgraphs). By symmetry, we may assume that $v(A)\geq v(B)$.
Case I: $v(B)\leq 2k$.
Assume that all vertices in $V(B)\setminus V(A)$ have degree greater than $k$. Let $b\in V(B)\setminus V(A)$. Since $|V(A)\cap V(B)|=k$, $b$ has a neighbor $c\in V(B)\setminus V(A)$. Since $B$ is triangle-free, the neighborhoods of $b$ and $c$ are disjoint subsets of $V(B)$. By the assumption, both contain at least $k+1$ vertices, which contradicts $v(B)\leq 2k$.
Hence, there is a vertex $b\in V(B)\setminus V(A)$ of degree at most $k$. By using the induction hypothesis for GbG-b,
$e(G)\penalty 10000\ =\penalty 10000\ e(G-b)+d_{G}(b)\penalty 10000\ \leq\penalty 10000\ e(G-b)+k\penalty 10000\ \leq\penalty 10000\ k\big(v(G)-1-k\big)+k\penalty 10000\ =\penalty 10000\ k\big(v(G)-k\big)\,.$
Case II: $v(B)>2k$.
Then also $v(A)>2k$, so we may use the induction hypothesis for both $A$ and $B$. Hence,
$e(G)\penalty 10000\ \leq\penalty 10000\ e(A)+e(B)\penalty 10000\ \leq\penalty 10000\ k\big(v(A)-k\big)+k\big(v(B)-k\big)\penalty 10000\ =\penalty 10000\ k\big(v(G)-k\big)\,.$
This proves the claim.
Hence, if $G$ is a triangle-free graph with average degree at least $2k$, then $v(G)\geq 2k$ and $e(G)\geq kv(G)>k\big(v(G)-k\big)$. Hence, $G$ contains a $(k+1)$-connected subgraph $G^{\prime}$. Since $G^{\prime}$ is also triangle-free, for some edge $vw\in E(G^{\prime})$, the neighborhoods of $v$ and $w$ are disjoint. This shows $v(G^{\prime})\geq d_{G^{\prime}}(v)+d_{G^{\prime}}(w)\geq 2(k+1)$. $\Box$
Already in [3], Mader proved the above result for graphs of girth larger than $k+1$. He also presented a construction of bipartite graphs of arbitrarily large girth and average degree arbitrarily close to $2k$ without a $(k+1)$-connected subgraph. Even the minimum degree version of Theorem On highly connected subgraphs is best possible, that is, there exists a bipartite graph with minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. One can build such a graph by gluing together copies of $H_{0}:=K_{k,2k-1}$ similarly to the final step of the proof of Theorem On highly connected subgraphs.
On highly edge-connected subgraphs
Mader proved in [4] that every (simple) graph $G$ with $e(G)>k\big(v(G)-\frac{k+1}{2}\big)$ and $v(G)\geq k$ contains a $(k+1)$-edge-connected subgraph. This condition is sharp for every $k$ and every $v(G)$: Consider the graph $G$ on a clique $A$ on $k$ vertices and an anticlique on $V(G)\setminus A$ with all possible edges between $A$ and $V(G)\setminus A$. Clearly, $G$ has not even a subgraph of minimum degree at least $k+1$. Note that, when $v(G)$ tends to $\infty$, the average degree of $G$ tends to $2k$.
We first show that an average degree of at least $2k$ has some stronger implications. For (simple) graphs in particular, it implies the existence of a $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
Theorem 4
Let $G$ be a loopless multigraph with e(G)>k(v(G)1)e(G)\,>\,k\,(v(G)-1). Then $G$ contains a submultigraph $G^{\prime}$ with
  1. (i)
    $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$.
  2. (ii)
    $G^{\prime}$ is $(k+1)$-edge-connected.
  3. (iii)
    $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees.
  4. (iv)
    If $G$ is a (simple) graph, then $v(G^{\prime})>2k$.
Proof. Let $G^{\prime}$ be a smallest nonempty induced submultigraph of $G$ with $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$. Let $\mathcal{P}$ be a partition of $V(G^{\prime})$ in more than one class. Then for all $A\in\mathcal{P}$, $e(G[A])\leq k(|A|-1)$. Hence the number of edges between distinct classes of $\mathcal{P}$ satisfies
[…]
$=\penalty 10000\ k\Big(v(G^{\prime})-1+|\mathcal{P}|-\sum_{A\in\mathcal{P}}|A|\Big)\penalty 10000\ =\penalty 10000\ k(|\mathcal{P}|-1)\,.$
In particular, it follows from the case $|\mathcal{P}|=2$, that $G^{\prime}$ is $(k+1)$-edge-connected. Moreover, by the Theorem by Tutte [5] and Nash-Williams [6], $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees. If $G$ and hence $G^{\prime}$ is a (simple) graph, then $G^{\prime}$ has more than $2k$ vertices, since otherwise $e(G^{\prime})\,\leq\,\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)\,\leq\,k\,(v(G^{\prime})-1)$. $\Box$
In particular, Theorem On highly edge-connected subgraphs applies to graphs of minimum degree at least $2k$. Even the minimum degree version is best possible in some senses:
  1. (1)
    The complete graph on $2k$ vertices has minimum degree 2k12k-1 but no $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
  2. (2)
    Every (2k1)(2k-1)-regular (simple) graph $G$ without a complete graph on $2k$ vertices as a component does not contain a subgraph $G^{\prime}$ with $k$ pairwise edge-disjoint spanning trees, because every subgraph $G^{\prime}$ has less than the necessary $k(v(G^{\prime})-1)$ edges:
    • If $v(G^{\prime})<2k$, then $e(G^{\prime})\leq\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)<k(v(G^{\prime})-1)$.
    • If $v(G^{\prime})=2k$, then $G^{\prime}$ is not complete, since otherwise it would be a component of $G$. Hence $e(G^{\prime})<\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)=k(v(G^{\prime})-1)$.
    • If $v(G^{\prime})>2k$, then $e(G^{\prime})\leq\frac{1}{2}(2k-1)v(G^{\prime})=kv(G^{\prime})-\frac{1}{2}v(G^{\prime})<k(v(G^{\prime})-1)$.
  3. (3)
    There exists a multigraph with minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph: Consider the graph $G$ on the vertex set $\{1,\dots,2(k+1)\}$ with k1k-1 parallel edges between each two consecutive vertices $i$ and $i+1$, one additional edge between $1$ and each of the $k$ vertices $\{2,\dots,k+1\}$, and between $2(k+1)$ and each of the $k$ vertices $\{k+2,\dots,2k+1\}$. Clearly, all vertices have degree 2k12k-1. Between the two vertex sets $\{1,\dots,k+1\}$ and $\{k+2,\dots,2k+2\}$ there are only k1k-1 edges. Hence, if $G^{\prime}$ is a $(k+1)$-edge-connected subgraph of $G$, then either $V(G^{\prime})\subseteq\{1,\dots,k+1\}$ or $V(G^{\prime})\subseteq\{k+2,\dots,2k+2\}$. In both cases, there is a vertex of degree at most $k$, $\max V(G^{\prime})$ or $\min V(G^{\prime})$, respectively. This is a contradiction.
  4. (4)
    A similar construction can be used for (simple) graphs. Together with ideas from Theorem On highly edge-connected subgraphs, one can show that for $k\leq 5$, a minimum degree $2k$ is still best possible: One can build graphs of minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph. Because the construction allows no generalization, we will not show it here. Instead, we now consider large $k$.
We now prove that, for (simple) graphs and for $k\to\infty$, already a minimum degree $\big(1+o(1)\big)k$ is sufficient for a $(k+1)$-edge-connected subgraph.
Let $\alpha\in\big(0,\frac{1}{2}\big)$ be fixed. All of the following definitions will depend on $k$, which will not be reflected in the notations. Let $m=m(k)\in[0,k]$ not necessarily an integer, which we will choose later.
Let $G$ be a graph and $v\in V(G)$ of degree $d_{G}(v)$. We define a delimited penalty
$m_{G}(v)\penalty 10000\ :=\penalty 10000\ \min\Big\{\max\big\{k+m-d_{G}(v),\,0\big\},\,m\Big\}\penalty 10000\ \in[0,m]\,.$
We consider the following weighted penalty of the graph $G$:
[…]
The sum of the first $l$ weights can be upper-bounded as follows:
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha j^{\alpha-1}dx\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ \int_{0}^{l}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ l^{\alpha}\,.$
Similarly, we obtain as a lower bound:
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha j^{\alpha-1}dx\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ \int_{1}^{l+1}\alpha x^{\alpha-1}dx\penalty 10000\ \geq\penalty 10000\ l^{\alpha}-1\,.$
Lemma On highly edge-connected subgraphsa
Let $G$ be a graph that contains a vertex $w$ of degree at most $k$. Then
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-w)-k^{\alpha}\,.$
Proof. Let $v_{1},\dots,v_{v(G)-1}$ be a list of $V(G-w)=V(G)\setminus\{w\}$ that attains
$M(G-w)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G-w}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)-1\}$, we have $d_{G}(v_{j})-d_{G-w}(v_{j})\in\{0,1\}$ as $G$ is simple. Since the function $x\mapsto\min\big\{\max\{k+m-x,\,0\},\,m\big\}$ is monotone decreasing and non-expanding, we have $m_{G-w}(v_{j})-m_{G}(v_{j})\leq d_{G}(v_{j})-d_{G-w}(v_{j})$. Hence $\Delta_{j}:=m_{G-w}(v_{j})-m_{G}(v_{j})\in[0,1]$ and
$\sum_{j=1}^{v(G)-1}\Delta_{j}\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{v(G)-1}d_{G}(v_{j})-d_{G-w}(v_{j})\penalty 10000\ =\penalty 10000\ d_{G}(w)\penalty 10000\ \leq\penalty 10000\ k\,.$
Consider the list $v_{1},\dots,v_{v(G)-1},w$ of the vertices of $G$. This yields
[…]
$>\penalty 10000\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\big(m_{G-w}(v_{j})-\Delta_{j}\big)\penalty 10000\ =\penalty 10000\ M(G-w)-\sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)-1}\Delta_{j}\leq k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, the weighted sum over the $\Delta_{j}\in[0,1]$ is maximized if $\Delta_{j}=1$ for $j\leq k$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-w)-\sum_{j=1}^{k}\alpha j^{\alpha-1}\penalty 10000\ \geq\penalty 10000\ M(G-w)-k^{\alpha}\,$,
using the upper bound on the sum of the first $k$ weights. $\Box$
Lemma On highly edge-connected subgraphsb
Let $G$ be a graph that contains a set $S$ of at most $k$ edges such that GSG-S is not connected, so it decomposes into non-empty subgraphs $A$ and $B$, that is, $V(G)$ is partitioned into $V(A)\cup V(B)$ and $E(G)\setminus S$ is partitioned into $E(A)\cup E(B)$. Then
$M(G)\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\,.$
Proof. Let $a_{1},\dots,a_{v(A)}$ and $b_{1},\dots,b_{v(B)}$ be lists of $V(A)$ and $V(B)$ that attain
$M(A)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(A)}\alpha j^{\alpha-1}m_{A}(a_{j})\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ M(B)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(B)}\alpha j^{\alpha-1}m_{B}(b_{j})\,.$
Clearly $d_{A}(a_{j})=d_{G-S}(a_{j})$ and hence $m_{A}(a_{j})=m_{G-S}(a_{j})$. Similarly $m_{B}(b_{j})=m_{G-S}(b_{j})$. We may assume by symmetry that $v(A)\leq v(B)$. By considering the list
$a_{1},b_{1},a_{2},b_{2},\dots,a_{v(A)},b_{v(A)},b_{v(A)+1},\dots,b_{v(B)}$
of all vertices of V(GS)=V(G)V(G-S)=V(G), we obtain
[…]
Let $v_{1},\dots,v_{v(G)}$ be a list of V(GS)=V(G)V(G-S)=V(G) that attains
$M(G-S)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G-S}(v_{j})\,.$
We have $d_{G}(v_{j})\geq d_{G-S}(v_{j})$ and hence $\Delta_{j}:=m_{G-S}(v_{j})-m_{G}(v_{j})\in[0,m]$ for every $j\in\{1,\dots,v(G)\}$. It is
$\sum_{j=1}^{v(G)}\Delta_{j}\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{v(G)}d_{G}(v_{j})-d_{G-S}(v_{j})\penalty 10000\ =\penalty 10000\ 2|S|\penalty 10000\ \leq\penalty 10000\ 2k\,.$
By considering the same list as a list of $V(G)$, we obtain
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G}(v_{j})\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\big(m_{G-S}(v_{j})-\Delta_{j}\big)\penalty 10000\ =\penalty 10000\ M(G-S)-\sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)}\Delta_{j}\leq 2k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, we obtain an upper bound of the weighted sum over the $\Delta_{j}\in[0,m]$ by setting $\Delta_{j}=m$ for $j\leq\big\lceil\frac{2k}{m}\big\rceil$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-S)-\sum_{j=1}^{\big\lceil\frac{2k}{m}\big\rceil}\alpha j^{\alpha-1}m\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\,$,
using the upper bound on the sum of the first $\big\lceil\frac{2k}{m}\big\rceil\leq\frac{4k}{m}$ weights. $\Box$
Theorem 5
For every $\alpha>0$, there is some $k_{0}$ such that, for all $k\geq k_{0}$, every graph with minimum degree at least $k+k^{\frac{1}{2}+\alpha}$ has a $(k+1)$-edge-connected subgraph.
Proof. Let $\alpha\in\big(0,\frac{1}{2}\big)$, $m=k^{\frac{1}{2}+\alpha}\leq k$ and $\mu=\frac{1}{2}m\sqrt{k}^{\alpha}$.
Let $G$ be a graph with $v(G)>k$ without a $(k+1)$-edge-connected subgraph. We prove by induction on $v(G)$ that $M(G)\geq\mu$.
A sequence of pairwise distinct vertices $w_{1},\dots,w_{l}\in V(G)$ is called detachable from $G$, if for each $j\in\{1,\dots,l\}$, $w_{j}$ has degree at most $k$ in $G-w_{1}-\dots-w_{j-1}$. We set
$z(G)\penalty 10000\ :=\penalty 10000\ \max\big\{l\penalty 10000\ :\penalty 10000\ w_{1},\dots,w_{l}\in V(G)\text{ is detachable from }G\big\}\,.$
Case I: $z(G)\geq\sqrt{k}$.
Then, there is a detachable sequence $w_{1},\dots,w_{\big\lceil\sqrt{k}\big\rceil}\in V(G)$. For each $j\in\{1,\dots,\big\lceil\sqrt{k}\big\rceil\}$$d_{G}(w_{j})\leq k+(j-1)$ and hence $m_{G}(w_{j})\geq m-(j-1)\geq m-(\big\lceil\sqrt{k}\big\rceil-1)\geq m-\sqrt{k}\geq\frac{2}{3}m$,for $k$ large enough. Using the lower bound on the sum of the first weights, this implies, again for $k$ large enough,
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{\big\lceil\sqrt{k}\big\rceil}\alpha j^{\alpha-1}m_{G}(v_{j})\penalty 10000\ \geq\penalty 10000\ \frac{2}{3}m\sum_{j=1}^{\big\lceil\sqrt{k}\big\rceil}\alpha j^{\alpha-1}\penalty 10000\ \geq\penalty 10000\ \frac{2}{3}m(\sqrt{k}^{\alpha}-1)\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}m\sqrt{k}^{\alpha}\penalty 10000\ =\penalty 10000\ \mu\,.$
Case II: $z(G)<\sqrt{k}$.
Then, there is a longest detachable sequence $w_{1},\dots,w_{z(G)}\in V(G)$. Hence, in the graph $G^{\prime}=G-w_{1}-\dots-w_{z(G)}$, all vertices have degree >k>k. Since $G^{\prime}$ is a non-empty subgraph of $G$, it is not $(k+1)$-edge-connected. Hence, there is a set $S$ of at most $k$ edges such that $G^{\prime}-S$ decomposes into two non-empty subgraphs $A$ and $B$. We may assume that all edges of $S$ are between $V(A)$ and $V(B)$.
Since $d_{G^{\prime}}(a)>k$ for all $a\in V(A)$, we have $v(A)>1$ and
$(v(A)-1)k\penalty 10000\ =\penalty 10000\ v(A)k-k\penalty 10000\ <\sum_{a\in V(A)}d_{G^{\prime}}(a)\penalty 10000\ -|S|\penalty 10000\ =\sum_{a\in V(A)}d_{A}(a)\penalty 10000\ \leq\penalty 10000\ v(A)(v(A)-1)\,$,
which yields $v(A)>k$, and similarly $v(B)>k$. Like $G$, also its subgraphs $A$ and $B$ do not contain a $(k+1)$-edge-connected subgraph, so we can apply the induction hypothesis to both $A$ and $B$. Together with Lemma On highly edge-connected subgraphsb, we obtain
$M(G^{\prime})\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\penalty 10000\ =\penalty 10000\ 2^{\alpha}\mu-m\Big(\frac{4k}{m}\Big)^{\alpha}\,.$
Finally we apply Lemma On highly edge-connected subgraphsa to get from $G$ to $G^{\prime}$ by removing the vertices $w_{1},\dots,w_{z(G)}$ one after another
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G^{\prime})-z(G)k^{\alpha}\penalty 10000\ \geq\penalty 10000\ 2^{\alpha}\mu-m\Big(\frac{4k}{m}\Big)^{\alpha}-k^{\frac{1}{2}}k^{\alpha}\,.$
For $k\to\infty$, both $m\big(\frac{4k}{m}\big)^{\alpha}=mo(k)^{\alpha}=o(\mu)$ and $k^{\frac{1}{2}}k^{\alpha}=o(m)k^{\alpha}=o(\mu)$. Hence, for $k$ large enough, the sum of both numbers is smaller than $(2^{\alpha}-1)\mu$, that is, $M(G)\geq\mu$.
This finishes the proof: If $G$ has minimum degree at least $k+m$, then $v(G)>k$ and $M(G)=0$, a contradiction. $\Box$
The condition $\alpha>0$ is sharp:
Theorem 6
For every constant $C$, there exists a $k\in\mathbb{N}$ and a (simple) graph with minimum degree at least $k+C\sqrt{k}$ without a $(k+1)$-edge-connected subgraph.
Proof. We may assume that $C\in\mathbb{N}$ and we choose $k$ as a square such that $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$. Further let $k$ be large enough such that $2^{2C^{2}}C\leq\sqrt{k}$.
We prove for $j=0,\dots,2C^{2}$ that there is a graph $G_{j}$ and $A_{j}\subseteq V(G_{j})$ with
  1. (i)
    $G_{j}$ does not contain a $(k+1)$-edge-connected subgraph.
  2. (ii)
    $|A_{j}|\penalty 10000\ =\penalty 10000\ 2^{j}(C\sqrt{k}+1)$.
  3. (iii)
    The vertices in $A_{j}$ have degree at least $k+\frac{1}{2C}j\sqrt{k}$.
  4. (iv)
    The vertices in $V(G_{j})\setminus A_{j}$ have degree at least $k+C\sqrt{k}$.
The claim follows with $j=2C^{2}$.
Define $G_{0}$ as a graph on $k+C\sqrt{k}+1$ vertices that contains an anticlique $A_{0}$ of $C\sqrt{k}+1$ vertices, a clique $V(G_{0})\setminus A_{0}$ on $k$ vertices and all possible edges between $A_{0}$ and $V(G_{0})\setminus A_{0}$. The vertices in $A_{0}$ have degree $k$ and those in $V(G_{0})\setminus A_{0}$ have degree $k+C\sqrt{k}$. A $(k+1)$-edge-connected subgraph of $G_{0}$ cannot contain any vertex from $A_{0}$, but also $G_{0}-A_{0}$ has too few vertices to contain a $(k+1)$-edge-connected subgraph.
Now let $0\leq j<2C^{2}$. Assume we have already constructed $G_{j},A_{j}$ and we want to construct $G_{j+1},A_{j+1}$. Note that
$|A_{j}|\penalty 10000\ =\penalty 10000\ 2^{j}(C\sqrt{k}+1)\penalty 10000\ \leq\penalty 10000\ 2^{j+1}C\sqrt{k}\penalty 10000\ \leq\penalty 10000\ 2^{2C^{2}}C\sqrt{k}\penalty 10000\ \leq\penalty 10000\ k\,.$
We build a graph $H_{j}$ from $G_{j}$ by adding an anticlique $B_{j}$ on $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$ vertices. From each vertex in $B_{j}$, we add one edge to every vertex in $A_{j}$ and $k-|A_{j}|$ more edges to other vertices from $G_{j}$. In $H_{j}$, the vertices in $B_{j}$ have degree $k$, the vertices in $A_{j}$ now have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ and all other vertices still have degree at least $k+C\sqrt{k}$. A $(k+1)$-edge-connected subgraph of $H_{j}$ cannot contain any vertex from $B_{j}$, but also $G_{j}=H_{j}-B_{j}$ does not contain one by the assumption.
We build $G_{j+1}$ from $H_{j}$ and one disjoint copy $H^{\prime}_{j}$. Let $A^{\prime}_{j}$ and $B^{\prime}_{j}$ be the copies of $A_{j}$ and $B_{j}$ in $H^{\prime}_{j}$. From every vertex in $B_{j}$, we add $C\sqrt{k}$ edges to vertices in $H^{\prime}_{j}$. Similarly, from every vertex in $B^{\prime}_{j}$, we add edges to vertices in $H_{j}$ until there are $C\sqrt{k}$ many. All together, we have added at most $2\frac{1}{2C}\sqrt{k}C\sqrt{k}=k$ edges between $H_{j}$ and $H^{\prime}_{j}$. Hence, a $(k+1)$-edge-connected subgraph of $G_{j+1}$ must be completely contained in either $H_{j}$ or $H^{\prime}_{j}$, which we prevented with the construction.
We set $A_{j+1}:=A_{j}\cup A^{\prime}_{j}$ of size $2|A_{j}|=2^{j+1}(C\sqrt{k}+1)$. The vertices in $A_{j+1}$ still have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ as in $H_{j}$. All other vertices, including those in $B_{j}\cup B^{\prime}_{j}$, have degree at least $k+C\sqrt{k}$. This finishes the construction. $\Box$
For the proof, we need $C=\Theta(\sqrt{\log k})$, which might be sharp: For some $D$, a minimum degree of at least $k+D\sqrt{k\log k}$ might be sufficient for a $(k+1)$-edge-connected subgraph.
Application on graph decomposition
The Theorems from the present paper have some applications to other issues. We want to mention decomposition problems of the following type, which were first studied by Thomassen [7]: What is the lowest number $k=k(s,t)$ such that every kk-(edge-)connected graph $G$ admits a partition of $V(G)$ into $A$ and $B$ such that $G[A]$ is ss-(edge-)connected and $G[B]$ is tt-(edge-)connected?
Corollary. Let $k_{1},k_{2}\geq 1$.
  1. (i)
    Let $G$ be a $(k_{1}+k_{2}+1)$-connected graph of minimum degree at least $3(k_{1}+k_{2})-1$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $|C_{i}|>2k_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-connected.
  2. (ii)
    Let $G$ be a triangle-free $(k_{1}+k_{2}+1)$-connected graph of minimum degree at least $2(k_{1}+k_{2})$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $G[C_{i}]$ is $(k_{i}+1)$-connected (and hence $|C_{i}|\geq 2(k+1)$).
  3. (iii)
    Let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of average degree at least $2(k_{1}+k_{2})+2$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $|C_{i}|>2k_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-edge-connected.
  4. (iv)
    There is a function $f(r)=\big(1+o(1)\big)r$ such that the following holds: Let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of minimum degree at least $f(k_{1}+k_{2})$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $G[C_{i}]$ is $(k_{i}+1)$-edge-connected.
Proof. The approach is due to Thomassen [7] and Stiebitz [8].
  • (i)
    $G$ has minimum degree at least $(3k_{1}-1)+(3k_{2}-1)+1$. By a result of Stiebitz [8], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $3k_{i}-1$. By Theorem On highly connected subgraphs, $G[A_{i}]$ has a $(k_{i}+1)$-connected subgraph on some vertex set $B_{i}$ of size $2k_{i}$, for $i\in\{1,2\}$. We choose maximal disjoint $C_{1},C_{2}\subseteq V(G)$ such that $C_{i}\supseteq B_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-connected, for $i\in\{1,2\}$.
    Assume that $X=V(G)\setminus(C_{1}\cup C_{2})\neq\emptyset$. Then $G[C_{1}\cup X]$ is not $(k_{1}+1)$-connected, that is, there is some $S_{1}\subseteq C_{1}\cup X$ of at most $k_{1}$ vertices such that $G[(C_{1}\cup X)\setminus S_{1}]$ has more than one connected component. One of them has to be disjoint from $C_{1}$. Let $Y\subseteq X$ be its vertex set.
    Now $G[C_{2}\cup Y]$ is not $(k_{2}+1)$-connected, that is, there is some $S_{2}\subseteq C_{2}\cup Y$ of at most $k_{2}$ vertices such that $G[(C_{2}\cup Y)\setminus S_{2}]$ has more than one connected component. One of them has to be disjoint from $C_{2}$. Let $Z\subseteq Y$ be its vertex set. The neighborhood of $Z$ is contained in $S_{1}\cup S_{2}$ of size at most $k_{1}+k_{2}$, a contradiction to the $(k_{1}+k_{2}+1)$-connectedness of $G$.
  • (ii)
    By a result of Kaneko [9] on triangle-free graphs, $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $2k_{i}$. Now use Theorem On highly connected subgraphs to obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-connected subgraphs. Then proceed as before.
  • (iii)
    By a result of Wang and Wu [10], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has average degree at least $2k_{i}$. Now use Theorem On highly edge-connected subgraphs to obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-connected subgraphs with vertex number greater than $2k_{1}$ and $2k_{2}$, respectively. Since Thomassen’s argument also works for edge-connectedness, we may proceed as before.
  • (iv)
    By Theorem On highly edge-connected subgraphs and Theorem On highly edge-connected subgraphs for small $k$, there exists a function $g(k)=\big(1+o(1)\big)k$$g(k)\leq 2k$ such that for every $k$, every graph of minimum degree at least $g(k)$ contains a $(k+1)$-edge-connected subgraph. We choose $f(r)=\max\big\{g(r-k)+g(k)+1\,\big|\,1\leq k\leq\frac{r}{2}\big\}$. We have that $f(r)=\big(1+o(1)\big)r$ for $r\to\infty$:
    If $k\geq\sqrt{r}$, then $g(r-k)+g(k)=\big(1+o(1)\big)(r-k)+\big(1+o(1)\big)k=\big(1+o(1)\big)r$.Otherwise, $g(r-k)+g(k)<\big(1+o(1)\big)(r-k)+2k\leq\big(1+o(1)\big)r+o(r)=\big(1+o(1)\big)r$.
    Now let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of minimum degree at least $f(k_{1}+k_{2})\geq g(k_{1})+g(k_{2})+1$. Again by Stiebitz [8], $V(G)$ has a partition into $A_{1}$ and $A_{2}$ such that, for $i\in\{1,2\}$, $G[A_{i}]$ has minimum degree at least $g(k_{i})$. We obtain disjoint $(k_{1}+1)$- and $(k_{2}+1)$-edge-connected subgraphs. Then proceed as before. $\Box$
The result of Wang and Wu [10] may also be applied to obtain average degree versions of (i) and (ii).